OpenStudy (anonymous):

Larry Klaus withdrew $310.80 simple interest from an account after 4 years. If he invested $1,110.00, how much was his rate of interest?

OpenStudy (anonymous):

OK whats the interest formula

OpenStudy (anonymous):

A=P(i+1)^n

OpenStudy (anonymous):

(1110+310.8)=1110(i+1)^4

OpenStudy (anonymous):

Now just solve for i!!!

OpenStudy (anonymous):

or dont, lol up to u